# Multi-scale Cascaded Large-Model for Whole-body ROI Segmentation

### v1

@@ Abstract
Organs-at-risk segmentation is critical for ensuring the safety and precision of radiotherapy and surgical procedures. However, existing methods for organs-at-risk image segmentation often suffer from uncertainties and biases in target selection, as well as insufficient model validation experiments, limiting their generality and reliability in practical applications. To address these issues, we propose an innovative cascaded network architecture called the Multi-scale Cascaded Fusing Network (MCFNet), which effectively captures complex multi-scale and multi-resolution features. MCFNet includes a Sharp Extraction Backbone and a Flexible Connection Backbone, which respectively enhance feature extraction in the downsampling and skip-connection stages. This design not only improves segmentation accuracy but also ensures computational efficiency, enabling precise detail capture even in low-resolution images. We conduct experiments using the A6000 GPU on diverse datasets from 671 patients, including 36,131 image-mask pairs across 10 different datasets. MCFNet demonstrates strong robustness, performing consistently well across 10 datasets. Additionally, MCFNet exhibits excellent generalizability, maintaining high accuracy in different clinical scenarios. We also introduce an adaptive loss aggregation strategy to further optimize the model training process, improving both segmentation accuracy and efficiency. Through extensive validation, MCFNet demonstrates superior performance compared to existing methods, providing more reliable image-guided support. Our solution aims to significantly improve the precision and safety of radiotherapy and surgical procedures, advancing personalized treatment. The code has been made available on GitHub: https://github.com/Henry991115/MCFNet.
Index Terms:
Whole-body ROI segmentation, multi-scale cascaded Large-Model, linear attention transformer, adaptive loss aggregation strategy.
@@ I Introduction
Medical image segmentation aims to quickly and accurately separate different tissues, organs, or lesion areas, facilitating clinical applications such as disease analysis, lesion detection, and treatment planning [1, 2]. In the field of medical image analysis, accurately identifying and segmenting malignant tumors is crucial for early cancer diagnosis, treatment planning, and efficacy evaluation. However, malignant tumors are often complex in shape, vary in size, have blurred boundaries, and are closely connected to surrounding tissues, posing significant challenges for accurate tumor identification and segmentation [3, 4, 5]. Accurately identifying and locating organs-at-risk (OAR) can help doctors and medical teams plan and execute treatment plans. In radiation therapy, precise identification of OAR can assist doctors in minimizing the impact of treatment on surrounding healthy tissues, thereby reducing adverse effects caused by the treatment [6, 7]. Tumor and OAR segmentation are fundamental to treatment planning in radiation therapy. Tumor segmentation helps doctors determine the target area for accurately aiming the radiation dose. Meanwhile, OAR segmentation allows doctors to identify and protect surrounding healthy tissues, avoiding unnecessary damage during treatment [8, 9].
@@
With the emergence of deep learning models, traditional methods constrained by human, financial, and time costs are gradually being abandoned [10, 11, 12, 13]. Deep learning-based models can automatically learn features from images and efficiently extract features of complex structures and textures, enabling the identification of ROI contours and facilitating subsequent medical image analysis [14, 15]. Most current efficient segmentation models require a substantial amount of high-performance GPU computing resources for training and have a strong dependency on large-scale annotated data. However, obtaining high-quality annotated data requires a significant amount of time and human resources [16, 17].
@@
Most current deep learning models for segmentation are based on UNet [18] and Transformer [19] architectures. UNet is renowned for its efficient and simple structure, which can effectively capture local features in images, while Transformers are known for their outstanding ability to capture global information [20, 21, 22]. The combination of UNet’s local feature extraction capabilities with the Transformer’s global information capture ability allows segmentation models to consider both local details and overall structure in images, thereby improving segmentation accuracy and robustness. This combination provides a more comprehensive and efficient solution for medical image analysis, offering strong support for advancements in the medical imaging field. However, the segmentation performance of most current methods is limited by the application scenario, and the robustness and generalization abilities of the models are usually weak. High-performance models like TransUNet [23], Swin-Unet [24], and MISSFormer [25] have demonstrated their capabilities primarily in segmenting eight organs in the heart and abdominal regions. UCTransNet [26] has expanded its tasks to include the segmentation of glands and cell nuclei. Nevertheless, in our actual experiments, we found that the segmentation performance of these methods varies significantly when applied to datasets with different image characteristics.
Existing works primarily focus on segmenting certain organs-at-risk and lesions, but they cannot segment multiple regions of the entire body, highlighting the limited application scenarios and the relatively poor robustness and generalization capabilities of these models. To address this, we propose conducting experiments on datasets that cover multiple organs-at-risk and tumors across the entire body. To enhance the model’s generalization ability for different image characteristics, we intend to capture image features from multiple angles using multi-scale and multi-resolution images, enabling the model to better adapt to current application scenarios. Finally, we design the FCB (Flexible Connection Backbone), which is targeted for multi-region segmentation, as one of the strong backbones. Additionally, we design a SEB (Sharp Extraction Backbone) to assist the FCB in better feature extraction. Cascading the two backbones can ensure high segmentation performance while designing the model to enhance its robustness and generalization ability. Our contributions are summarized as follows:
@@
    Novel Multi-scale Cascaded Large-Model: We propose a novel Multi-scale Cascaded U-shaped Network for whole-body ROI segmentation. This network cascades two backbone networks, namely our designed FCB (Flexible Connection Backbone) and SEB (Sharp Extraction Backbone). The two backbones are used to handle images of different scales and resolutions, with feature fusion occurring in the four skip connections and the down-sampling bottom layers of the network.
@@
    Innovative Adaptive Loss Aggregation Strategy: We introduce a novel adaptive loss aggregation strategy named Adaptive-MFA (Adaptive Multi-scale Feature-Mixing Loss Aggregation). By using this strategy, we can create new prediction maps by combining available prediction maps, thereby optimizing the model training as much as possible.
@@
    Extensive Validation for Robustness and Generalization: To thoroughly validate the model’s robustness and generalization capabilities, we selected ten datasets targeting organs-at-risk and tumors throughout the body for experiments. The segmented regions include the brain, thoracic cavity, abdomen, and lower limbs, with two of these datasets provided by our collaborating hospital.
@@ II Related Work
In this section, we introduce commonly used architectures for medical image segmentation as well as cascade network architectures.
@@ II-A Medical Image Segmentation
Medical image segmentation refers to accurately separating different structures (such as organs, lesion regions, etc.) in medical images, which is a crucial task in medical image analysis and computer-aided diagnosis. To address this task, research on U-shaped networks based on encoder-decoder structures has seen explosive growth. UNet innovatively introduced the encoder-decoder architecture, enabling the aggregation of features from different stages through skip connections. Subsequently, UNet++ [28] and UNet 3+ [29] improved the skip connection mechanism, further enhancing the effectiveness of feature aggregation.
In recent years, the application of Transformers in medical image segmentation has attracted widespread attention. Originally used in the field of natural language processing, Transformers have shown great potential in computer vision tasks due to their powerful modeling capabilities and flexibility. Compared to traditional convolutional neural networks (CNNs), Transformer architectures primarily rely on self-attention mechanisms, allowing them to capture long-range dependencies in images globally. ViT [30] is one of the first models to apply Transformers to image classification tasks, processing images by splitting them into fixed-size patches and treating these patches as sequences inputted to the Transformer for processing. The same concept has been introduced into medical image segmentation, such as the TransUNet model, which significantly improves segmentation performance by combining the local feature extraction capability of UNet with the global contextual modeling capability of Transformers. Swin Transformer is a hierarchical Transformer architecture that further enhances computational efficiency and feature extraction capability through a sliding window mechanism. Swin-Unet applies this architecture to medical image segmentation tasks and achieves excellent results. Recently, Rahman et al. proposed the Cascaded MERIT [31] model, which effectively improves the accuracy and stability of medical image segmentation through techniques such as multiscale hierarchical feature extraction and cascaded attention decoding.
@@ II-B Cascaded Networks for Medical Image Segmentation
In recent years, cascade networks have gained considerable attention in medical image segmentation. These networks concatenate multiple modules, each handling specific subtasks, gradually improving overall segmentation performance. Cascade networks can integrate feature information at different scales and decompose tasks into multiple stages for optimization, with each stage utilizing feature representations from the preceding stage [31]. Information propagation is achieved through mechanisms such as skip connections and attention mechanisms. In experimental validation, cascade networks have demonstrated effectiveness and superiority in medical image segmentation tasks, offering significant potential as diagnostic and therapeutic tools in clinical medicine.
Recently, the CASCADE model introduced a novel attention-based decoder that learns local (contextual) relationships between pixels, addressing the shortcomings of Transformer models in learning local relationships. G-CASCADE [32] refines multi-stage feature maps generated by a hierarchical Transformer encoder through graph convolutional blocks, contributing to improved accuracy and robustness of segmentation results through progressive refinement. The Cascaded MERIT [31] model improves generalization capability by introducing a multiscale self-attention mechanism. Traditional single-scale self-attention mechanisms may not effectively capture multiscale information in images, whereas Cascaded MERIT computes self-attention at multiple scales to comprehensively understand image features, thereby enhancing model performance. MCANet proposes a multi-scale cross-axis attention mechanism that efficiently captures information from different sizes and shapes of lesion regions or organs in medical images. By computing dual cross attentions between two parallel axial attentions, better global information capture is achieved, facilitating accurate segmentation of target areas in medical images.
@@ III Proposed Method
In this section, we introduce the multi-scale cascade network model MCFNet designed for whole-body ROI segmentation. First, we provide an overall introduction to the model, then we describe each component of the model in detail, and finally, we explain the adaptive loss aggregation strategy we designed.
Figure 1: Illustration of the overall architecture of MCFNet.
III-A MCFNet for Whole-body ROI Segmentation
Fig. 1 shows the overall architecture of our designed MCFNet network. MCFNet consists of three parts: Cascaded Backbones, Decoders with cascaded skip connections, and Aggregation. The Cascaded Backbones module is composed of SEB (Acute Extraction Backbone) and FCB (Elegant Connection Backbone), enabling multi-scale and multi-resolution input for the model. The Decoders with cascaded skip connections module mainly achieves feature aggregation at skip connections and bridge layers in the cascade network, enhancing feature representation by fully utilizing feature information from different levels and different sub-networks. The Aggregation module primarily reflects the aggregation of the decoder output layers and the application of our designed adaptive loss aggregation strategy.
@@ III-B Cascaded Backbones in MCFNet
The cascaded backbone network consists of SEB (Sharp Extraction Backbone) and FCB (Flexible Connection Backbone). These two sub-backbone networks are responsible for feature extraction from images of different resolutions. Generally, high-resolution images can provide richer details and information, aiding the model in more accurately identifying and segmenting structures and regions within the images. However, this increases computational complexity and resource requirements. Low-resolution images are typically more robust to noise and interference but lose detailed information. Considering application requirements and computational resources, we choose to use FCB for processing $224\times 224$ images and SEB for assisting with the processing of $256\times 256$ images. In the encoder stage, the FCB sub-network uses traditional max pooling and convolution operations. This means that FCB has some limitations in feature extraction during the downsampling stage. Therefore, the design of SEB focuses on feature extraction during the downsampling stage while ensuring a relatively small number of parameters and computational load.
Specifically, SEB enhances the recognition and extraction of detailed features during the downsampling stage for high-resolution images by adding SE attention mechanisms to the encoding layers. As the feature extraction process deepens in SEB, the images go through BatchNorm, MaxPool, and convolution operations via the SEB Stem and encoding layers. The prediction map of SEB is finally output from the highest layer of the SEB decoder and is subjected to a Hadamard product operation with the input image of SEB to obtain a new output image. This new output image is then resized to the $224\times 224$ input image required by FCB. The low-resolution images are input into FCB and processed through the FCB Stem and encoding layers. The four layers of feature maps are output through a skip connection mechanism and subsequently aggregated with those from SEB.
@@ III-C Cascaded Skip-connection Module in MCFNet
As shown in Fig. 2, this section mainly introduces the cascaded skip-connection mechanism. The efficient backbone model, FCB, is used to ensure the effectiveness of feature extraction, while a lightweight and simple SEB is employed to capture detailed information in high-resolution images, thus compensating for the limitations of FCB.
Figure 2: (a) Illustration of the CSM (Cascaded Skip-connection Module). (b) Illustration of the LAT (Linear Attention Transformer Block) in the CSM.
Specifically, when the encoding layers of FCB reach the lowest level after downsampling, they enter the CAB (Cross-Attention Bridge Layer). This module compensates for the loss of low-level features and then aggregates features with the output feature maps of the fourth encoding layer in SEB. For skip connections in FCB, the features first pass through the LAT (Linear Attention Transformer Block), enabling comprehensive and efficient feature extraction across both channel and spatial dimensions. As shown in Fig. 2, the structure of LAT is based on our published work, FSCA-Net, specifically the PAT (Parallel Attention Transformer Block). We make new improvements to accelerate the convergence speed of the transformer-like structure and reduce the number of model parameters caused by its architecture. After the input tensor enters the LAT module, we first use a linear layer for dimensionality reduction to align the tensor dimensions with the channel numbers. The compressed dimensional feature maps are then processed sequentially through layers including LN, 2D-SCCA attention, and another LN layer. Finally, we obtain the output features by passing the data through a linear layer for dimensional restoration.
Subsequently, the feature maps outputted by LAT are aggregated with the feature maps outputted by the skip connections of resized SEB. The aggregated feature maps are then fed into the decoding layer of FCB through skip connections, which directly propagate shallow features to deeper layers, preserving fine-grained information captured by shallow networks. In SEB, the output prediction map ($S_{output}$) from the highest decoding layer is fed into the SAM (Sigmoid Activation Map Module). This module mainly utilizes a 1x1 convolution to output a single-channel feature map, and then applies the $Sigmoid$ activation function to nonlinearly reweight features, enhancing the representation of key features while suppressing unimportant ones. The calculation formula for the SAM module is as follows:
@@
Figure 3: Illustration of the Aggregation module.
where the target image coordinates are denoted as $(x,y)$, the original image coordinates are $(i,j)$, and the four nearest neighbor pixels in the original image are $(i_{1},j_{1})$, $(i_{2},j_{2})$, $(i_{3},j_{3})$, $(i_{4},j_{4})$. Additionally, $\alpha=x-i_{1}$, $\beta=y-j_{1}$, and $pixel(i,j)$ represent the pixel value at coordinates $(i,j)$ in the original image. Each decoding layer of SEB and FCB generates a prediction map, representing the predicted results for the input. These prediction maps come from different scales and levels, reflecting different levels of feature information and semantic understanding. We feed these prediction maps into the Aggregation module for the aggregation of multi-scale information, aiming to comprehensively utilize information from different scales and improve the accuracy and robustness of the final prediction results.

### v2

# Multi-scale Cascaded Foundation Model for Whole-body Organs-at-risk Segmentation

@@ Abstract
Accurate segmentation of organs-at-risk (OARs) is vital for safe and precise radiotherapy and surgery. Most existing studies segment only a limited set of organs or regions, lacking a systematic treatment of OARs segmentation. We present a Multi-scale Cascaded Fusion Network (MCFNet) that aggregates features across multiple scales and resolutions. MCFNet consists of a Sharp Extraction Backbone for the downsampling path and a Flexible Connection Backbone for skip-connection fusion, strengthening representation learning in both stages. This design improves boundary localization and preserves fine structures while maintaining computational efficiency, enabling reliable performance even on low-resolution inputs. Experiments on an NVIDIA A6000 GPU using 36,131 image–mask pairs from 671 patients across 10 datasets show consistent robustness and strong cross-dataset generalization. An adaptive loss-aggregation strategy further stabilizes optimization and yields additional gains in accuracy and training efficiency. Through extensive validation, MCFNet outperforms existing methods, excelling in organ segmentation and providing reliable image-guided support for computer-aided diagnosis. Our solution aims to improve the precision and safety of radiotherapy and surgery while supporting personalized treatment, advancing modern medical technology. The code has been made available on GitHub: https://github.com/Henry991115/MCFNet.
@@ I Introduction
Medical image segmentation aims to accurately separate different tissues, organs, or lesion areas, supporting clinical tasks like disease analysis, lesion detection, and treatment planning [1, 2]. Accurate segmentation of malignant tumors is crucial for early cancer diagnosis, treatment planning, and efficacy evaluation. However, tumors often have complex shapes, blurred boundaries, and are closely connected to surrounding tissues, posing challenges for segmentation [3, 4, 5, 6]. Identifying and locating OARs is vital in radiation therapy, helping minimize damage to healthy tissues and improving treatment outcomes [7, 8, 9]. Tumor and OAR segmentation are essential for precise treatment planning, enabling targeted radiation while protecting healthy tissues [10, 11].
Figure 1: MCFNet is trained on ten diverse datasets to perform whole-body organs-at-risk (OARs) segmentation, covering the head & neck, thorax, abdomen, prostate, and femur regions.
@@
With the rise of deep learning, traditional methods constrained by human and financial costs are being replaced [12, 13, 14, 15]. Deep learning models can automatically extract complex features from images, enabling efficient OARs contour identification and further medical analysis [16, 17, 18]. However, most current segmentation models require substantial GPU resources for training and large-scale annotated datasets, which are time-consuming and resource-intensive to create [19, 20].
@@
Most current deep learning segmentation models are based on UNet [21] and Transformer [22] architectures. UNet is known for its simple structure that effectively captures local features, while Transformers excel at capturing global information [23, 24, 25]. Combining UNet’s local feature extraction with the Transformer’s global information capture improves segmentation accuracy and robustness by considering both local details and the overall structure in images. This combination enhances medical image analysis, supporting progress in the field [26, 27, 28]. However, the segmentation performance of existing methods is often limited by the application scenario, with weak robustness and generalization abilities. High-performance models such as TransUNet [29], Swin-Unet [30], and MISSFormer [31] perform well in segmenting organs in the heart and abdominal regions, while UCTransNet [32] can segment glands and cell nuclei. Nevertheless, in our actual experiments, we find that the segmentation performance of these methods varies significantly when applied to datasets with different image characteristics [33].
In this paper, we propose MCFNet, a multi-scale cascaded fusion framework for robust segmentation of OARs and lesions under diverse imaging protocols and anatomical regions. MCFNet synchronizes multi-scale and multi-resolution inputs through two complementary branches: the Flexible Connection Backbone (FCB) for coarse global semantics and the Sharp Extraction Backbone (SEB) for fine boundary details. A Linear Attention Transformer is embedded in FCB to model long-range dependencies efficiently, while SEB focuses on high-frequency structures. An adaptive multi-scale loss aggregation strategy further enhances accuracy and training stability. As shown in Fig. 1, we evaluate MCFNet on ten datasets spanning head-neck, thorax, abdomen, prostate, and femur regions. Extensive experiments and ablation studies demonstrate its strong robustness, generalization, and parameter efficiency.
The main contributions of our work are as follows:
@@
    We propose MCFNet, a novel multi-scale cascaded network that integrates the SEB and FCB to encode and fuse features across scales and resolutions. Feature fusion is performed at four skip connections and the bottleneck layers, enabling both boundary fidelity and semantic consistency. Comprehensive evaluations on 10 heterogeneous datasets show that the proposed framework is robust, computationally efficient, and generalizes well.
@@
    We develop a Linear Attention Transformer Block (LAT) within the FCB to enhance feature extraction along the channel and spatial axes. LAT reduces quadratic complexity to linear, thereby accelerating convergence and improving scalability. This design enables the FCB to capture global dependencies with fewer parameters, achieving a favorable accuracy–efficiency trade-off.
@@
    We propose an Adaptive Multi-scale Feature-Mixing Loss Aggregation (Adaptive-MFA) strategy, which dynamically generates new prediction maps by combining multi-scale outputs. This adaptive aggregation stabilizes training and further improves segmentation performance.
The structure of this study is as follows: Section II discusses related work. The methodology of our proposed model is detailed in Section III. The experiment configuration is introduced in Section IV. Section IV describes the results and discussion, and the effectiveness of the proposed method is validated. Finally, Section V presents the summarization of our research conclusions.
Figure 2: Discussion on Model Complexity on the CPCGEA Dataset. (a) Visualization of model performance and parameter size. (b) Visualization of model performance and FLOPs.
@@ II Related Work
This section introduces common architectures for medical image segmentation and cascade networks.
@@ II-A Medical Image Segmentation
Medical image segmentation, which involves accurately separating different structures like organs and lesions, is essential for medical image analysis and diagnosis. The introduction of encoder-decoder-based U-shaped networks, such as UNet, has significantly advanced this task, with subsequent improvements like UNet++ [34] and UNet 3+ [35] enhancing skip connection mechanisms for better feature aggregation.
In recent years, Transformers have gained attention in medical image segmentation. Although initially designed for NLP, Transformers excel in computer vision tasks due to their powerful modeling and flexibility [36]. Unlike CNNs, Transformer models employ self-attention to model long-range dependencies on a global scale. ViT [37] introduce Transformers to image classification, using fixed-size patches to segment images, treating them as sequences for processing. This idea has been adapted to medical image segmentation, exemplified by the TransUNet model, which combines UNet’s local feature extraction with the global context modeling of Transformers. The Swin Transformer, a hierarchical model with sliding windows, further boosts feature extraction and efficiency. The Cascaded MERIT model by Rahman et al. [38] has been shown to improve segmentation accuracy and stability using multiscale features and cascaded attention decoding.
Figure 3: Illustration of the overall MCFNet architecture. The network is built upon two complementary backbones: the Sharp Extraction Backbone for fine details and the Flexible Connection Backbone for global semantics, which are integrated via a Cascaded Skip-connection Module and an Aggregation Module.
@@ II-B Cascaded Networks for Medical Image Segmentation
In recent years, cascade networks have become prominent in medical image segmentation. These networks concatenate multiple modules to handle specific subtasks, improving segmentation performance gradually. By integrating feature information at different scales and decomposing tasks into multiple stages, cascade networks optimize segmentation through mechanisms like skip connections and attention mechanisms [39]. Experimental results show that cascade networks are effective and often superior for medical image segmentation, with consistent gains in Dice and HD95 across diverse benchmarks (e.g., cardiac structures, and tumor/OARs delineation)[40]. Beyond accuracy, cascaded designs improve robustness to domain shifts and small, low-contrast targets, translating into more reliable auto-contouring and reduced manual correction time in clinical workflows—thereby offering tangible benefits for diagnostics, radiotherapy planning, and longitudinal treatment assessment[41, 42].
The CASCADE model introduces an attention-based decoder that captures local pixel-wise dependencies, addressing a known limitation of Transformer architectures. G-CASCADE [43] employs graph-convolution blocks to refine hierarchical Transformer encoder features, improving segmentation accuracy and robustness. The Cascaded MERIT [38] model improves generalization by employing a multiscale self-attention mechanism, which better captures multiscale information than traditional single-scale attention mechanisms. MCANet [44] employs multi-scale cross-axis attention to model lesions and organs with heterogeneous sizes and shapes, strengthening global context and enabling precise target segmentation.
Distinguished from existing studies, our method offers enhanced multi-scale modeling, demonstrating leading performance and generalization in robust multi-dataset evaluations.
Figure 4: (a) Illustration of the Cascaded Skip-connection Module (CSM). (b) Illustration of the Linear Attention Transformer Block (LAT) in the CSM.
@@ III Proposed Method
In this section, we introduce the multi-scale cascade network model MCFNet designed for whole-body OARs segmentation. First, we provide an overall introduction to the model, then we describe each component of the model in detail, and finally, we explain the adaptive loss aggregation strategy we designed.
III-A MCFNet for Whole-body OARs Segmentation
Fig. 3 shows the overall architecture of our designed MCFNet network. MCFNet consists of three parts: Cascaded Backbones, Decoders with cascaded skip connections, and Aggregation. The Cascaded Backbones module is composed of Sharp Extraction Backbone (SEB) and Flexible Connection Backbone (FCB), enabling multi-scale and multi-resolution input for the model. The Decoders with cascaded skip connections module mainly achieves feature aggregation at skip connections and bridge layers in the cascaded network. The Aggregation module primarily reflects the aggregation of the decoder output layers and the application of our designed adaptive loss aggregation strategy.
@@ III-B Cascaded Backbones in MCFNet
The cascaded backbone network consists of SEB and FCB, responsible for extracting features from images of different resolutions. High-resolution images provide rich details but increase computational complexity. Low-resolution images are more robust to noise but lose details. We use FCB for $224\times 224$ images and SEB for $256\times 256$ images. FCB uses max pooling and convolution in the encoder, but the downsampling stage can further enhance feature capture. SEB focuses on improving feature extraction during downsampling with fewer parameters and computational load.
Specifically, SEB enhances the extraction of detailed features during downsampling for high-resolution images by incorporating SE attention mechanisms in the encoding layers. The images pass through BatchNorm, MaxPool, and convolution operations in the SEB Stem and encoding layers. The final prediction map is produced from the SEB decoder’s highest layer, and a Hadamard product is applied with the input image to generate a new output image. This output is resized to the $224\times 224$ input required by FCB. Low-resolution images are processed in the FCB through its Stem and encoding layers, with feature maps output via skip-connections and subsequently aggregated with those from the SEB.
@@ III-C Cascaded Skip-connection Module in MCFNet
As shown in Fig. 4, this section introduces the cascaded skip-connection mechanism. The efficient FCB backbone ensures effective feature extraction, while the lightweight SEB captures detailed information in high-resolution images to complement FCB’s limitations.
Specifically, when the encoding layers of FCB reach the lowest level after downsampling, they enter the Cross-Attention Bridge Layer proposed in our previous work, which compensates for the loss of low-level features and aggregates them with the output from SEB’s fourth encoding layer. FCB’s skip connections first pass through the Linear Attention Transformer Block (LAT), which enables efficient feature extraction across both channel and spatial dimensions. The LAT architecture is based on the parallel attention transformer module (PAT) from our previous work FSCA-Net, and has been improved to accelerate convergence and reduce the number of parameters. After the input tensor passes through the LAT module, it first undergoes dimensionality reduction via a linear layer to align the tensor dimensions with the channel numbers, reducing computational and storage demands. The compressed feature maps then pass through LN, 2D-SCCA attention, and another LN layer, ultimately being restored to their original dimensions through a linear layer.
The feature maps from LAT are aggregated with those from resized SEB and passed into FCB’s decoding layer via skip connections, preserving shallow feature details. In SEB, the highest decoding layer’s output ($S_{output}$) is processed by the SAM module, which uses 1x1 convolution and the $Sigmoid$ activation function to enhance key features and suppress less important ones. The calculation formula for the SAM module is as follows:
@@
where the target image coordinates are denoted as $(x,y)$, the original image coordinates are $(i,j)$, and the four nearest neighbor pixels in the original image are $(i_{1},j_{1})$, $(i_{2},j_{2})$, $(i_{3},j_{3})$, $(i_{4},j_{4})$. Additionally, $\alpha=x-i_{1}$, $\beta=y-j_{1}$, and $pixel(i,j)$ represent the pixel value at coordinates $(i,j)$ in the original image. Each decoding layer of SEB and FCB generates a prediction map at different scales, reflecting varying feature levels. These maps are input into the Aggregation module to combine multi-scale information, enhancing the accuracy and robustness of the final predictions.
Figure 5: Illustration of the Aggregation module that merges hierarchical multi-scale features from FCB and SEB into a compact representation supervised by Adaptive-MFA.